OpenStudy (anonymous):

I am stuck on this one problem, can someone please help me? in 1993 the life expectancy of males in a certain counrty was 63.2. In 1997, it was 65.5 years. Let E represent the life expectancy in year T and let T represent the number of years since 1993.

OpenStudy (anonymous):

so

OpenStudy (anonymous):

wat do u want to find

OpenStudy (anonymous):

some how I keep getting 57.5, but this does not look right...

OpenStudy (chaise):

Let: (x,y) = (E,T) (0, 63.2) (4, 65.5) Do you understand how I got past the first step? In order to find the gradient of your line, based on previous years. m = (y1-y2)/(x1-x2) m= (E1-E2)/(T1-T2) m = (63.2-65.5)/(0-4) m = -2.3/-4 m = 23/40 The equation: y = mx+c can be used, you already know your gradient which is 23/40, and you know a plot on your graph (T,E) y = mx+c E = mT+c 65.5 = (23/40)x-4+c 65.5 =-2.3+c c = 67.8 Your equation is: E = 23/40T + 67.8 This is not right, but I cannot find my error. Could it be that I have my x and y's round the wrong way? I didn't think it mattered. I hope you're asking for an equation of the data.

OpenStudy (anonymous):

so would the life expectancy in 2007 67.8? I am not sure, this is why I am asking. but the way you displayed does make some sense

OpenStudy (chaise):

Oh crap, I stuffed the whole thing up, sigh. I am much better working with pen to paper.

OpenStudy (chaise):

Let me try again.

OpenStudy (anonymous):

I have done this for the past hour and keep coming up with something different each time.. This should be simple for me, but not this time.. I am sorry

OpenStudy (chaise):

It's fine, let me do it with pen to paper :) Make things easier

OpenStudy (anonymous):

Thank you for all your help, I really appreciate it :)

OpenStudy (chaise):

Okay, got it. Your equation is: E=mT+c Where E is the life expectancy and T is the time after 1993 and m = 23/40 and c = 63.2 Amatuer mistake, I'm ashamed. So to determine the life expectancy at 2007 (and 14 years after 1993) E = mT+c E = (23/40)x14 + 63.2 E = 8.05 +63.2 E = 71.25 Sorry I took so long. If you need help on how I got the equation, I'll be happy to go through it with you. Your E is your Y, Your T is your X m = (y1-y2)/(x1-x2) m = (e1-e2)/(t1-t2) m = (0-4)/(63.2-65.5_ m = 23/40 I'm sure you know how to get your equation from this point onwards :)

OpenStudy (anonymous):

This is what I came up with this last time too. It is ok, I am not much help on this end. You have been such a great help, I am just sorry I have been such a pain with this one... I guess after working it over and over for more than an hour can make you a little crazy (so to speak)

OpenStudy (chaise):

Hehe, not a problem. I should have known, just got my x and y axis mixed around. I should have known. Sorry :P Anything else I can do for you? I don't mind helping with you genuinely appreciate the help.

OpenStudy (anonymous):

That was the last one I needed help with, but thank you again